You want to buy a new car for $74,300, and the finance office at the dealership has quoted you a loan with an APR of 6.7% for 72 months to buy the car. What will your monthly payment be? What is the effective annual rate on this loan? Please show all answers and calculations.
Answer #1

Annuity Payment = r(PV)/[1 - (1 + r)-n]

Annuity Payment = (0.067/12)(74,300)/[1 - (1 + 0.067/12)-72]

Annuity Payment = $1,256.07

So,

Monthly Payment = $1,256.07

EAR = (1 + 0.067/12)12 - 1

EAR = 6.91%

Please "Like" if you find this helps and comment for your doubts.